OpenStudy (anonymous):

The typical healthy resting heart rate in adults is 60–80 beats per minute. If you live 80 years, how many times will your heart beat?

OpenStudy (mathstudent55):

Ok. This is just a long multiplication. How many minutes in an hour?

OpenStudy (anonymous):

60

OpenStudy (mathstudent55):

How many hours in a day?

OpenStudy (anonymous):

24

OpenStudy (mathstudent55):

How many days in year?

OpenStudy (mathstudent55):

365.25

OpenStudy (mathstudent55):

\(80 ~years \times \dfrac{365.25~days}{year} \times \dfrac{24~hours}{day} \times \dfrac{60~minutes}{hour} \times \dfrac{60~beats}{minute}\)

OpenStudy (anonymous):

I got 252,460,800

OpenStudy (mathstudent55):

Multiply it out with 60 at the end as written above. Then do it also with 80 at the end. Then you can answer the range is from _____ to _____. Alternatively, you can calculate it once with 70 at the end.

OpenStudy (mathstudent55):

I get 2,524,608,000

OpenStudy (anonymous):

I must of did something wrong

OpenStudy (mathstudent55):

Your result is 10 times smaller than it would be. Maybe you entered one of the 60's as just 6.

OpenStudy (mathstudent55):

60 beats per minute: 2,524,608,000 80 beats per minute: 3,366,144,000 You can answer approx. 2.5 billion to 3.4 billion Using the average, 70 beats per minute: 2,945,376,000 Approx. 2.9 billion

OpenStudy (anonymous):

So how do i put that into one of these answers because these are the choices a)3 \[10^{6}\] b)3 \[10^{15}\] c)3 \[10^{12}\] d)3 \[10^{9}\]

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Ok. Now you need the average value as a number in scientific notation.

OpenStudy (mathstudent55):

We get an average value of 2,945,376,000. You can round this off to 3,000,000,000 (3 billion).

OpenStudy (mathstudent55):

3,000,000,000 = 3 * 1,000,000,000

OpenStudy (mathstudent55):

1 = 10^0 10 = 10^1 100 = 10^2 1000 = 10^3 10,000 = 10^4 etc. 1,000,000,000 = 10^9

OpenStudy (mathstudent55):

3,000,000,000 = 3 * 1,000,000,000 = 3 * 10^9

OpenStudy (anonymous):

So is it D

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

Yay!!!!

OpenStudy (anonymous):

Thank you

OpenStudy (mathstudent55):

yw